# vApps: Verifiable Applications at Internet Scale

### v1

@@ Abstract.
Blockchain technology promises a decentralized, trustless, and interoperable infrastructure. However, widespread adoption remains hindered by issues such as limited scalability, high transaction costs, and the complexity of maintaining coherent verification logic across different blockchain layers. This paper introduces Verifiable Applications (vApps), a novel development framework designed to streamline the creation and deployment of verifiable blockchain computing applications. vApps offer a unified Rust-based Domain-Specific Language (DSL) within a comprehensive SDK, featuring modular abstractions for verification, proof generation, and inter-chain connectivity. This eases the developer’s burden in securing diverse software components, allowing them to focus on application logic. The DSL also ensures that applications can automatically take advantage of specialized precompiles and hardware acceleration to achieve consistently high performance with minimal developer effort, as demonstrated by benchmark results for zero-knowledge virtual machines (zkVMs). Experiments show that native Rust execution eliminates interpretation overhead, delivering up to an 832x cycle count improvement compared to EVM-based approaches. Precompiled circuits accelerate proving by over 95%, while GPU acceleration boosts throughput by up to 30x and recursion compresses proof size by up to 230x, enabling succinct and efficient verification. The framework also supports seamless integration with Web2 and Web3 systems, enabling developers to focus solely on their application logic. Through modular architecture, robust security guarantees, and composability, vApps pave the way toward a trust-minimized and verifiable Internet-scale application environment.
@@ 2.1. Rollups
As shown in Figure 2(b), modular blockchain architecture separates the execution layer from the settlement layer. Users interact with application logic on the L2 layer, where a Prover generates an integrity proof for each batch of transactions. This proof is submitted to the settlement layer to verify correctness and challenge invalid or censored transactions. The settlement layer provides security for proof verification, data commitment, and state finality. Offloading execution to L2 reduces the L1’s computational burden, addressing the scalability limits of monolithic designs. The modular architecture can extend execution layers beyond simple transaction processing. For instance, emerging environments such as the vApp SDK can support general-purpose computation or data processing, going beyond typical rollup use cases.
@@ 2.2. Verifiable Computing
Verifiable computing is a model in which a Prover executes a function $y=F(x,w)$ and produces a succinct proof attesting that the resulting output $y$ is correct, given a program $F$, public inputs $x$, and a $w$, as shown in Figure 1. A Verifier then checks this succinct proof, confirming it without having to re-run the function. Verifiable computing techniques rely on various trust assumptions and may use software-based or hardware-based attestation. The three types of verifiable computing techniques are:
  • (1) Optimistic: This technique optimistically executes transactions off-chain and posts state updates without immediate proofs. Watchers can challenge potentially fraudulent transactions by submitting fraud proofs during a defined period. On the settlement layer, fraud proofs are verified, and invalid updates can be reverted. The trust assumption is that there is at least 1 honest and incentivized watcher who will monitor the rollup and submit fraud proofs when necessary. The complexity of the code used to verify fraud proofs introduces a significant risk, as its large and intricate nature increases the likelihood of bugs.
  • (2) Zero Knowledge (ZK) Proofs: This technique uses cryptographic (zk-SNARK (snark, ) or zk-STARK (stark, )) methods to generate succinct off-chain attestations of correct computations that produce state updates. These succinct proofs are then efficiently verified on-chain without dispute. The trust assumption is in the soundness of the underlying ZK proofs and the correctness of their implementation. Recent advancements in zero knowledge virtual machines (zkVMs) have made it possible to integrate ZK proof verification in rollup and decentralized application development at scale. Section 6 provides a deeper technical exposition of this promising approach.
@@
Rollup solutions, built on previously discussed verifiable computing techniques, support diverse application requirements according to each technique’s performance and security trade-offs, as shown in Figure 3.
@@ 6.3. Performance
We map the individual steps in proving to stages S0 to S4 from Figure 8. This mapping demonstrates that Ethereum execution follows the standard vApp transaction lifecycle. We see that the zkVM spends a large fraction of its cycle count (approaching $60\%$ in some cases), in the block execution phase, corresponding to S2, the state transition function. We note that recover_senders is optimized via an ECDSA recovery precompile. In Table 5, we show that executing programs inside the EVM can cause up to 832x cycle count overhead relative to simply executing native Rust in the zkVM. Combined with the transaction lifecycle profiling for Reth in Table 2, this allows us to realize opportunities for dramatic optimization. For instance, the block_execution phase can be optimized with native Rust, and initialize_witness_db, which sets up an authenticated data structure for the account data and storage required for transaction execution, can be optimized with precompiles and specialized data structures (qmdb, ). Further, deserialize_inputs can be optimized with a ZK friendly serializer.
@@
deserialize_inputs (S0) 15.25 % 17.40 % 23.52 %
recover_senders (S0) 2.62 % 3.22 % 6.21 %
block_validation (S0 and S4) 1.32 % 1.33 % 1.79 %
initialize_witness_db (S1) 8.20 % 9.31 % 14.64 %
block_execution (S2 and S4) 59.63 % 53.84 % 31.58 %
state_root_computation (S3) 11.68 % 13.51 % 20.36 %
accrue_logs_bloom (S4) 0.63 % 0.61 % 0.71 %
other (S0 and S4) 0.68 % 0.79 % 1.18 %
Table 2. Cycle counts by transaction lifecycle steps
@@ 6.3.4. Repricing the EVM for zkVMs
We illustrate this in Table 5, where we show cycle counts and the overhead over native Rust for various programs in SP1. We compare Rust execution on three representative Fibonacci programs to Wasmi, which is a Web Assembly interpreter, and Revm, which is an interpreter for the EVM. Our results demonstrate that the EVM can cause up to 832x overhead even for benign programs like Fibonacci. This overhead occurs as a result of executing EVM-specific opcodes inside the zkVM. In the vApp framework, developers can write their applications directly in Rust and execute the program natively in the zkVM, avoiding interpreter overhead altogether. The code to reproduce these experiments is available at
@@ 6.3.6. Parallelization: Sharding and Recursion
zkVMs are subject to even greater parallelization if proving is run on a cluster of many GPUs. Modern GPUs, with thousands of cores and high memory bandwidth, efficiently handle cryptographic operations such as Fast Fourier Transforms (FFTs) specific to ZK proving. Properly optimized zkVM implementations demonstrate near-linear scaling in throughput with additional GPUs. The ubiquitous availability of GPU clusters due to datacenter buildouts in related industries such as AI generates complementarities for zkVMs. Increasingly, production workloads for ZK rollups are being proven in clusters of hundreds of GPUs and novel marketplaces for coordinating proving providers are emerging.
@@ Acknowledgements.
We gratefully acknowledge the valuable feedback and assistance provided by individuals and teams regarding system design and benchmarking. We also extend our appreciation to Patrick O’Grady of Commonware for his insightful suggestions that improved the manuscript.

### v2

@@ Abstract.
Blockchain technology promises a decentralized, trustless, and interoperable infrastructure. However, widespread adoption remains hindered by issues such as limited scalability, high transaction costs, and the complexity of maintaining coherent verification logic across different blockchain layers. This paper introduces Verifiable Applications (vApps), a novel development framework designed to streamline the creation and deployment of verifiable blockchain computing applications. vApps offer a unified Rust-based Domain-Specific Language (DSL) within a comprehensive SDK, featuring modular abstractions for verification, proof generation, and inter-chain connectivity. This eases the developer’s burden in securing diverse software components, allowing them to focus on application logic. The DSL also ensures that applications can automatically take advantage of specialized precompiles and hardware acceleration to achieve consistently high performance with minimal developer effort, as demonstrated by benchmark results for zero-knowledge virtual machines (zkVMs). Experiments show that native Rust execution eliminates interpretation overhead, delivering up to an 832x cycle count improvement compared to EVM-based approaches. Precompiled circuits can accelerate the proof by more than 95%, while GPU acceleration increases throughput by up to 30x and recursion compresses the proof size by up to 230x, enabling succinct and efficient verification. The framework also supports seamless integration with the Web2 and Web3 systems, enabling developers to focus solely on their application logic. Through modular architecture, robust security guarantees, and composability, vApps pave the way toward a trust-minimized and verifiable Internet-scale application environment.
@@ 2.1. Rollups
As shown in Figure 2(b), modular blockchain architecture separates the execution layer from the settlement layer. Users interact with application logic on the L2 layer, where a Prover generates an integrity proof for each batch of transactions. This proof is submitted to the settlement layer to verify correctness and challenge invalid or censored transactions. The settlement layer provides security for proof verification, data commitment, and state finality. Offloading execution to the L2 reduces the L1’s computational burden, addressing the scalability limits of monolithic designs. The modular architecture can extend execution layers beyond simple transaction processing. For instance, emerging environments such as the vApp SDK can support general-purpose computation or data processing, going beyond typical rollup use cases.
@@ 2.2. Verifiable Computing
Verifiable computing is a model in which a Prover executes a function $y=F(x,w)$ and produces a succinct proof attesting that the resulting output $y$ is correct, given a program $F$, public inputs $x$, and a witness $w$, as shown in Figure 1. A Verifier then checks this succinct proof, confirming it without having to re-run the function. Verifiable computing techniques rely on various trust assumptions and may use software-based or hardware-based attestation. The three types of verifiable computing techniques are:
  • (1) Optimistic: This technique optimistically executes transactions off-chain and posts state updates without immediate proofs. Watchers can challenge potentially fraudulent transactions by submitting fraud proofs during a defined period. On the settlement layer, fraud proofs are verified, and invalid updates can be reverted. The trust assumption is that there is at least 1 honest and incentivized watcher who will monitor the rollup and submit fraud proofs when necessary. The complexity of the code used to verify fraud proofs introduces a significant risk, and its large and intricate nature increases the likelihood of bugs.
  • (2) Zero Knowledge (ZK) Proofs: This technique uses cryptographic (zk-SNARK (snark, ) or zk-STARK (stark, )) methods to generate succinct off-chain attestations of correct computations that produce state updates. These succinct proofs are then efficiently verified on-chain without dispute. The trust assumption is in the soundness of the underlying ZK proofs and the correctness of their implementation. Recent advancements in zero knowledge virtual machines (zkVMs) have made it possible to integrate ZK proof verification in rollup and decentralized application development at scale. Section 6 provides a deeper technical exposition and empirical study of this promising approach.
@@
Rollup solutions, built on the previously discussed verifiable computing techniques, support diverse application requirements according to each technique’s performance and security trade-offs, as shown in Figure 3.
@@ 6.3. Performance
We map the individual steps in proving Ethereum blocks to stages S0 to S4 from Figure 8. This mapping demonstrates that Ethereum execution follows the standard vApp transaction lifecycle. We see that the zkVM spends a large fraction of its cycle count (approaching upwards of $63.33\%$ in some cases), in the block execution phase, corresponding to S2, the state transition function. For small blocks, only 28.28% of cycles are spent on execution, as the cost of data access is not well amortized across a small number of transactions. We note that recover_senders is optimized via an ECDSA recovery precompile.
In Table 5, we show that executing programs inside the EVM can cause up to 832x cycle count overhead relative to simply executing native Rust in the zkVM. Combined with the transaction lifecycle profiling for Reth in Table 2, this allows us to realize opportunities for dramatic optimization. For instance, the block_execution phase can be optimized with native Rust, and initialize_witness_db, which sets up an authenticated data structure for the account data and storage required for transaction execution, can be optimized with precompiles and specialized data structures (qmdb, ). Further, deserialize_inputs can be optimized with a ZK friendly serializer.
@@
deserialize_inputs (S0) 13.26 % 17.81 % 24.07 %
recover_senders (S0) 1.97 % 2.86 % 5.50 %
block_validation (S0 and S4) 1.18 % 1.41 % 1.92 %
initialize_witness_db (S1) 13.48 % 18.05 % 24.08 %
block_execution (S2 and S4) 63.33 % 50.47 % 28.28 %
state_root_computation (S3) 5.82 % 8.23 % 14.57 %
accrue_logs_bloom (S4) 0.57 % 0.64 % 0.75 %
other (S0 and S4) 0.38 % 0.53 % 0.83 %
Table 2. Cycle counts by transaction lifecycle steps.
@@ 6.3.4. Repricing the EVM for zkVMs
We illustrate this in Table 5, where we show cycle counts and the overhead over native Rust for various programs in SP1. We compare Rust execution on three representative Fibonacci programs to Wasmi, which is a Web Assembly interpreter, and Revm, which is an interpreter for the EVM. Our results demonstrate that the EVM can cause up to 832x overhead even for benign programs like Fibonacci. This overhead occurs as a result of executing EVM-specific opcodes inside the zkVM. In the vApp framework, developers can write their applications directly in Rust and execute the program natively in the zkVM, avoiding interpreter overhead altogether. It also highlights the potential of rearchitecting the EVM using RISC-V to accelerate proof generation. The code to reproduce these experiments is available at
@@ 6.3.6. Parallelization: Sharding and Recursion
zkVMs are subject to even greater parallelization if proving is run on a cluster of many GPUs. Modern GPUs, with thousands of cores and high memory bandwidth, efficiently handle cryptographic operations such as Fast Fourier Transforms (FFTs) specific to ZK proving. Properly optimized zkVM implementations demonstrate near-linear scaling in throughput with additional GPUs. The ubiquitous availability of GPU clusters due to datacenter buildouts in related industries such as AI generates complementarities for zkVMs. Increasingly, production workloads for ZK rollups are being proven in clusters of thousands of GPUs and novel marketplaces for coordinating proving providers are emerging.
@@ Acknowledgements.
We gratefully acknowledge the valuable feedback and assistance provided by individuals and teams regarding system design and benchmarking. We also extend our appreciation to Patrick O’Grady of Commonware, Ye Zhang of Scroll and Tarun Chitra of Robot Ventures for their insightful suggestions that improved the manuscript.